OpenStudy (anonymous):

How many moles of H2 would be contained in 4.0 L of the gas at 202.6 kPa and 127°C? A. 89.6 mol B. 6.38 mol C. 0.24 mol D. 0.77 mol

OpenStudy (anonymous):

@JFraser i think its (B)

OpenStudy (jfraser):

use the ideal gas law\[P*V = n*R*T\]just be sure to convert your temp to kelvin, just like before

OpenStudy (jfraser):

that's not what i get, check your values and make sure you're using the proper value for R based on the units of pressure

OpenStudy (anonymous):

i got (D)

OpenStudy (jfraser):

if you got D, then you forgot to convert the temperature to kelvin

OpenStudy (anonymous):

oh. then let me redo it

OpenStudy (anonymous):

A

OpenStudy (jfraser):

it's not A, either. let me show you the setup

OpenStudy (jfraser):

starting with\[P*V = n*R*T\]is rearranged to solve for \(n\)\[n = \frac{P*V}{R*T}\]

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

or was i right with b?

OpenStudy (jfraser):

\[n = \frac{(202.6kPa)*(2.0L)}{(8.31\frac{kPa*L}{mol*K})*(400K)}\]

OpenStudy (jfraser):

the answer is C)

OpenStudy (jfraser):

when you work with gases, the temp has to be in K